OpenStudy (anonymous):

Solve system by graphing y=2x+4 y=-x+7

OpenStudy (anonymous):

Graph both lines on the same coordinate system. The intersection point of the two lines is the solution of the system.

OpenStudy (anonymous):

set them equal and solve \[\huge 2x+4=-x+7\] in three steps

OpenStudy (anonymous):

Still dont get it

OpenStudy (anonymous):

satellite73, the question asks for a graphical solution.

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

a solution is a solution if you can't solve \[2x+4=-x+7\] then there the problem is bigger than graphing

OpenStudy (usukidoll):

dude just assign random values of x and make a table. Then graph the equation. ONe with blue one with red and then where it intersects there is the solution. What is this SO? You speak of?

OpenStudy (usukidoll):

*throws a canteen at satellite*

OpenStudy (usukidoll):

I know what the answer is...just not saying it until the graphs appear :)